1. 4. How many grams of glucose are present in 0.5L of 2.0M glucose solution. The molecular mass of glucose is 180 g/mol.
2. How many milliliters of a 1.00 % KCl solution are needed to provide a patient with 5.00 g KCI?
1
Expert's answer
2020-09-23T04:35:49-0400

1-

Molarity "= \frac{no of moles}{volume of solution in liters}"

"=" "\frac{wt of glucose}{0.5}"


"= 2" mol/lit

"= \frac{x}{(180\times0.5)}"

"= 180" grams of glucose

2-

"1%"% Kcl is 1 gram in 100 gm

so 5 gms in 500 gm

water density is 1gm/cc

so 500gm = 500 ml or 0.5liter